### v1

@@ 5.1 Experimental details
Implementation. We formulate the task as binary classification, where label 0 denotes real videos and label 1 denotes AI-generated videos. We train the model using the AdamW optimizer [adamw] with a learning rate of 3e-4 and weight decay of 0.05. A cosine learning rate schedule is applied with 1 warmup epoch and a minimum learning rate of 1e-6. We set the maximum gradient norm to 1.0 and enable automatic mixed precision (AMP) for efficient training. Each experiment is trained with 3 seeds. For the base AE-Swin model, we use a hidden dimension of 512 with 8 attention heads, and V-JEPA2 [assran2025vjepa2selfsupervisedvideo] as the vision encoder. The temporal and spatial window sizes are set to 4, with 2 transformer blocks for both temporal and spatial modeling. In the base setting, each video is sampled to produce 16 embeddings, and the same frames are used across all benchmarks and ablations for fairness. Although V-JEPA2 processes 32 raw frames due to its 2-frame tubelet design, it still outputs 16 embeddings, ensuring a consistent number of input tokens. All experiments are conducted on a single NVIDIA RTX 6000 Ada GPU with 48GB memory.
@@
Benchmark & Metrics. We benchmark our method against recent SoTA approaches published in top-tier venues or widely adopted as strong baselines including DeMamba [demamba], NPR [npr24], STIL [stil], TALL [tall], ResTraV [interno2025aigenerated_restrav], D3 [Zheng_2025_D3], WaveRep Augmentation [corvi2025waverep], Forgelens [Chen_2025_forgelen], Effort [yan2025orthogonal_effort], and NSG-VD [zhang2025physicsdriven_nsgvd]. When training code is publicly available, we re-implement the methods using our dataset; otherwise, we perform inference using the authors’ released pretrained weights (WaveRep Augmentation [corvi2025waverep]). For evaluation, we report Accuracy, Precision, Recall, F1-score, and AUC to comprehensively assess classification performance, robustness, and discrimination capability across different video generators.
5.2 Experimental results
@@
Table 1 reports the benchmark results on the seen generators. Traditional methods such as D3 struggle severely, performing close to random guessing (approximately 0.51 accuracy), highlighting the difficulty of modern high-quality video generators. More recent models including ResTraV, NPR, STIL, TALL, WaveRep, and DeMamba show progressively stronger performance, with DeMamba reaching 0.9515 average accuracy. In contrast, EA-Swin significantly outperforms all baselines, achieving 0.9866 average accuracy and 0.9991 AUC, demonstrating both near-perfect discrimination capability and stable performance across diverse commercial generators. The consistent improvement across all metrics (Accuracy, Precision, Recall, F1-score, AUC) confirms the effectiveness of our spatiotemporal modeling design.
Table 2 presents results on unseen generators to evaluate cross-distribution generalization. While several prior methods experience noticeable performance degradation (particularly WaveRep and TALL on certain generators), EA-Swin demonstrates robustness, achieving $0.974$ average accuracy and 0.997 AUC. Notably, the model preserves high Precision and Recall across nearly all unseen generators, indicating that it captures intrinsic generative artifacts rather than overfitting to specific training distributions. These results demonstrate that EA-Swin not only achieves state-of-the-art performance on seen data but also generalizes effectively to emerging and previously unseen video generation models.
Among the baselines, we observe that embedding–based trajectory methods (ResTraV & D3) show clear limitations. Specifically, ResTraV suffers from an information bottleneck due to aggressive dimensionality reduction, while D3 collapses to near-random performance by predicting almost all videos as fake (recall equals 1), revealing the weakness of heuristic variance-based criteria under modern high-quality generators. Although WaveRep Augmentation achieves competitive results, it mainly provides a data augmentation strategy; with a large-scale dataset such as EA-Video, its relative advantage becomes less pronounced. Among the baselines, DeMamba performs the best and confirms the importance of structured spatiotemporal modeling, yet it relies on a relatively large and computationally heavy architecture. Several other models exhibit noticeable performance drops on unseen generators, suggesting limited robustness and reliance on generator-specific artifacts.
6 Ablation Study
@@
Vision encoder. We evaluate the impact of different ViT-based vision backbones on our framework, including V-JEPA2 [assran2025vjepa2selfsupervisedvideo], CLIP [CLIP], DINOv3 [simeoni2025dinov3], and DINOv2 [dinov2]. As shown in Table 3, V-JEPA2 consistently achieves the best performance on both validation and test sets, attaining the highest Accuracy, F1-score, and AUC, while CLIP remains competitive but slightly inferior. DINOv3 and DINOv2 show noticeably lower results, particularly on the test set where DINOv2 suffers the largest drop, indicating weaker generalization. Overall, these results suggest that stronger self-supervised spatiotemporal representations, as learned by V-JEPA2, provide more discriminative features for AI-generated video detection.
@@
Metric Val set Test set
VJEPA2 CLIP DINO3 DINO2 VJEPA2 CLIP DINO3 DINO2
[…]
AUC 0.999 0.999 0.995 0.990 0.997 0.997 0.970 0.957
@@
Number of input frames. We further investigate the impact of reducing the number of input frames on model performance. As shown in Figure 9, decreasing the frame count from 16 to 8, 4, and 2 leads to a gradual decline across all metrics, with Recall and F1 being more sensitive to frame reduction. Nevertheless, the performance drop remains moderate, indicating that EA-Swin maintains reasonable robustness even under limited temporal information, while a higher number of frames still provides more stable and discriminative representations.
@@ 7 Conclusion
We presented EA-Swin, an enhanced Swin Transformer framework that improves feature representation and contextual modeling for robust visual recognition. By integrating the proposed enhancement modules into the hierarchical transformer backbone, EA-Swin effectively strengthens multiscale feature aggregation and long range dependency modeling while maintaining computational efficiency. Extensive experiments demonstrate that the proposed approach consistently outperforms baseline models across multiple benchmarks, validating its effectiveness and generalization capability. These results confirm that carefully designed architectural refinements can further unlock the potential of transformer-based vision models.
However, several limitations remain. First, the additional enhancement components introduce moderate computational and memory overhead, which may limit deployment in highly resource-constrained environments. Second, the current evaluation focuses primarily on standard benchmarks, and broader testing in real-world or domain-specific scenarios is still needed to fully assess robustness. Future work will explore lightweight design strategies to reduce complexity, adaptive mechanisms to dynamically allocate attention across scales, and broader applications such as dense prediction and cross-modal tasks. Additionally, investigating self-supervised or foundation-model pretraining strategies may further improve scalability and performance.
@@ Acknowledgements
We would like to express our gratitude to our colleagues at N2TP who assisted in the research. The primary author would like to thank Quang Hung Nguyen (Viettel) for his assistance in data collection.
@@ 0.A.2 More AI-generated video detection methods
Multimodal large language models (MLLMs). Another line of work explores the use of MLLMs for AI-generated video detection: BusterX [wen2025busterxmllmpoweredaigeneratedvideo, wen2026busterxunifiedcrossmodalaigenerated], Skyra [li2025skyraaigeneratedvideodetection], Vidguard-R1 [park2025vidguardr1aigeneratedvideodetection], $M$-Det [song2024_mm_det], DeepTraceReward [fu2025learninghumanperceivedfakenessaigenerated], AIGVE [xiang2025aigvetoolaigeneratedvideoevaluation]. While these approaches benefit from strong semantic understanding and interpretability, they suffer from two key limitations. As MLLMs are typically large and highly general-purpose, making them computationally expensive and poorly suited for scalable video-level detection. Moreover, several studies indicate that such models often focus on describing video content rather than performing true forensic detection, effectively assessing whether the model can reason about or narrate potential artifacts instead of learning discriminative signals for real-versus-generated classification. As a result, MLLM-based approaches remain more aligned with video understanding or analysis tasks than robust, standalone video detection.
@@ 0.B.2 Video frames from generators
Appendix 0.C Extended results
Below, we present more experiment result. First, the hyperparameters sweep is shown in Tables 4 & 5. Second, we present the result of ablation study experimented on each vision encoders: V-JEPA 2 (Tables 6 & 7), CLIP (Tables 9 & 9), DINO-v3 (Tables 10 & 11), DINO-v2 (Tables 12 & 13).

### v2

@@ 5.1 Experimental details
Implementation. We train a binary classifier (0: real, 1: AI-generated) using AdamW with learning rate $3\times 10^{-4}$, weight decay $0.05$, cosine decay with 1 warmup epoch and minimum learning rate $10^{-6}$. Gradients are clipped to $1.0$ and automatic mixed precision (AMP) is enabled. All experiments are run with 3 random seeds on a single NVIDIA RTX 6000 Ada GPU (48GB).
The base AE-Swin uses a hidden size of $512$, 88 attention heads, and V-JEPA2 as the vision encoder. Temporal and spatial window sizes are both set to $4$ with two transformer blocks for each. Every video produces $16$ embeddings; V-JEPA2 consumes $32$ frames via tubelets but outputs $16$ tokens, ensuring consistent inputs across experiments. More details of configuration are given in the Supplementary Material.
Baselines & Metrics. We benchmark our method against recent SoTA approaches published in top-tier venues or widely adopted as strong baselines including DeMamba [demamba], NPR [npr24], STIL [stil], TALL [tall], ResTraV [interno2025aigenerated_restrav], D3 [Zheng_2025_D3], WaveRep Augmentation [corvi2025waverep], Forgelens [Chen_2025_forgelen], Effort [yan2025orthogonal_effort], and NSG-VD [zhang2025physicsdriven_nsgvd]. We re-implement all the methods using our dataset; only for WaveRep Augmentation, we used the pretrained weight since the training code is not released. For evaluation, we report Accuracy, Recall, F1-score, and AUC to comprehensively assess classification performance, robustness, and discrimination capability across different video generators.
5.2 Results
Table 1 reports the benchmark results on the seen generators. Traditional methods such as D3 struggle severely, performing close to random guessing (approximately 0.51 accuracy), highlighting the difficulty posed by modern high-quality video generators. More recent models including ResTraV, NPR, STIL, TALL, WaveRep, and DeMamba show progressively stronger performance, with DeMamba reaching 0.9515 average accuracy. Recently proposed detectors such as Forgelens and NSG-VD exhibit mixed results, with Forgelens achieving strong performance (0.977 accuracy) while NSG-VD remains unstable across generators. In contrast, EA-Swin consistently outperforms all baselines, achieving 0.9866 average accuracy, 0.9869 F1, and 0.9991 AUC, demonstrating both near-perfect discrimination capability and stable performance across diverse commercial generators. The consistent improvement across all metrics confirms the effectiveness of our spatiotemporal modeling design.
@@ 5.2 Results
Table 2 presents results on unseen generators to evaluate cross-distribution generalization. Several prior methods experience noticeable performance degradation, most prominently WaveRep, which collapses on SKR, PV, Pika2, and Gen4 (e.g., 0.503/0.539/0.418/0.389 accuracy), and TALL, which also drops substantially on these generators. DeMamba remains the strongest baseline with an average accuracy of 0.922 and 0.948 AUC, while Forgelens shows high overall scores (0.882 accuracy, 0.971 AUC) but still exhibits instability on challenging cases such as Gen4. In contrast, EA-Swin demonstrates robust generalization, achieving 0.974 average accuracy and 0.997 AUC, while maintaining high Recall (Avg Recall 0.965) across nearly all unseen generators. These results suggest that EA-Swin captures more transferable generative artifacts rather than overfitting to the training distribution, delivering SoTA performance on both seen and emerging video generation models.
@@
Among the baselines, we observe that the embedding-based statistical methods (ResTraV, D3, NSG-VD) show clear limitations. Specifically, ResTraV and NSG-VD suffer from an information bottleneck due to aggressive dimensionality reduction, while D3 collapses to near-random performance by predicting almost all videos as fake (recall equals 1), revealing the weakness of heuristic variance-based criteria under modern high-quality generators. Although WaveRep Augmentation achieves competitive results, it mainly provides a data augmentation strategy; with a large-scale dataset such as EA-Video, its relative advantage becomes less pronounced. Among the baselines, DeMamba performs the best and confirms the importance of structured spatiotemporal modeling, yet it relies on a relatively large and computationally heavy architecture. Several other models exhibit noticeable performance drops on unseen generators, suggesting limited robustness and reliance on generator-specific artifacts.
5.3 Ablation Study
@@
Vision encoder. We evaluate the impact of different ViT-based vision backbones on our framework, including V-JEPA2 [assran2025vjepa2selfsupervisedvideo], CLIP [CLIP], DINOv3 [simeoni2025dinov3], and DINOv2 [dinov2], and a ViT-like encoder: ConvNeXt-v2 [convnextv2]. As shown in Table 3, V-JEPA2 consistently achieves the best performance on both validation and test sets, attaining the highest Accuracy, F1-score, and AUC, while CLIP remains competitive but slightly inferior. DINOv3 and DINOv2 show noticeably lower results, particularly on the test set where DINOv2 suffers the largest drop, indicating weaker generalization. Overall, these results suggest that stronger self-supervised spatiotemporal representations, as learned by V-JEPA2, provide more discriminative features for AI-generated video detection.
@@
Backbone Val set Test set
Acc Prec Recall F1 AUC Acc Prec Recall F1 AUC
VJEPA2 0.986 0.986 0.987 0.987 0.999 0.975 0.985 0.966 0.975 0.997
CLIP 0.987 0.983 0.991 0.987 0.999 0.974 0.983 0.965 0.974 0.997
DINO3 0.971 0.976 0.968 0.972 0.995 0.891 0.970 0.811 0.865 0.970
DINO2 0.954 0.956 0.953 0.955 0.990 0.874 0.949 0.791 0.846 0.957
ConvNeXt2 0.977 0.980 0.975 0.978 0.997 0.916 0.976 0.858 0.913 0.976
@@
5.4 Robustness Test
Number of input frames. We show the impact of reducing the number of input frames on model performance. As shown in Figure 10, decreasing the frame count from 16 to 8, 4, and 2 leads to a gradual decline across all metrics, with Recall and F1 being more sensitive to frame reduction. Nevertheless, the performance drop remains moderate, indicating that EA-Swin maintains reasonable robustness even under limited temporal information, while a higher number of frames still provides more stable and discriminative representations.
Table 4: Robustness test for EA-Swin on the validation set
Base Blur Comp. Noise
Acc 0.974 0.955 0.931 0.916
Prec 0.983 0.969 0.976 0.994
Recall 0.965 0.942 0.938 0.841
F1 0.974 0.955 0.956 0.912
AUC 0.997 0.991 0.990 0.988
Robustness Evaluation. To evaluate robustness to common real-world video post-processing, we generate three validation variants using ffmpeg: H.264 re-encoding for compression (CRF 36), Gaussian noise with optional downscaling and temporal–uniform noise injection (CRF 40), and Gaussian blur ($\sigma=2$). Such perturbations commonly appear in videos shared on social media platforms due to re-encoding, resizing, and transmission artifacts. As shown in Table 4, EA-Swin maintains stable performance across all perturbations with only moderate degradation from the clean setting (Acc 0.974, AUC 0.997). Under blur and compression, accuracy remains above 0.93 and AUC stays around 0.99, indicating strong resilience to realistic re-encoding artifacts. Gaussian noise is the most challenging condition, where accuracy drops to 0.916 and recall to 0.841, yet AUC remains high at 0.988. Overall, these results demonstrate consistent robustness to common video degradations.
@@ 6 Conclusion
We presented EA-Swin, an embedding-agnostic spatiotemporal detection framework for AI-generated video detection. Our results demonstrate that modeling the dynamics of pretrained video representations provides strong and consistent improvements in detection over prior pixel-level and trajectory-based approaches. These findings suggest that modern AI-generated video detection should shift from pixel-space analysis toward representation-space modeling, where temporal consistency and higher-level structure remain difficult for generative models to reproduce.
More broadly, this work highlights the growing importance of representation-level forensics in the era of foundation video models. As generative systems continue to improve visual realism, detection methods must increasingly rely on higher-level spatiotemporal signals rather than visible artifacts. We hope that EA-Swin and the EA-Video benchmark will serve as a foundation for future research on scalable and robust synthetic video detection.
@@ Acknowledgements
We would like to express our gratitude to our colleagues at N2TP (Phong Ho, Nhung Duong, Trang Pham) for their assistance with the research. The primary author would like to thank Quang Hung Nguyen (Viettel) for his assistance in data collection.
@@ 0.A.2 More AI-generated video detection methods
Multimodal large language models (MLLMs). Another line of work explores the use of MLLMs for AI-generated video detection: BusterX [wen2025busterxmllmpoweredaigeneratedvideo, wen2026busterxunifiedcrossmodalaigenerated], Skyra [li2025skyraaigeneratedvideodetection], Vidguard-R1 [park2025vidguardr1aigeneratedvideodetection], MM-Det [song2024_mm_det], DeepTraceReward [fu2025learninghumanperceivedfakenessaigenerated], AIGVE [xiang2025aigvetoolaigeneratedvideoevaluation]. While these approaches benefit from strong semantic understanding and interpretability, they suffer from two key limitations. As MLLMs are typically large and highly general-purpose, making them computationally expensive and poorly suited for scalable video-level detection. Moreover, several studies indicate that such models often focus on describing video content rather than performing true forensic detection, effectively assessing whether the model can reason about or narrate potential artifacts instead of learning discriminative signals for real-versus-generated classification. As a result, MLLM-based approaches remain more aligned with video understanding or analysis tasks than robust, standalone video detection.
@@ 0.B.1 Dataset detail
We provide detailed statistics of the dataset composition in Table 5. The table reports the number of AI-generated videos per generator and split, along with the corresponding number of paired real videos used as negative samples. As discussed, the training and test splits share the same generator families, while the validation split is constructed from unseen generators to evaluate out-of-distribution generalization.
Table 5: Dataset composition per generator and split
Split Generator #AI_vids %in_split #correspond_real Published
Train set
train veo3 5054 13.691 - 7/2025
train sora2 4857 13.158 - 9/2025
train cogvideox 4605 12.475 - 3/2025
train hunyuan 4524 12.256 - 3/2025
train easyanimate 4199 11.375 - 7/2024
train ltxvideo 4130 11.188 - 12/2024
train pika1 3388 9.178 - 12/2023
train wan2 2155 5.838 - 4/2025
train kling2 2056 5.570 - 4/2025
train sora 1917 5.193 - 2/2024
Val set
val veo3 2213 13.989 2119 -
val sora2 2104 13.300 2016 -
val hunyuan 1969 12.446 1886 -
val ltxvideo 1870 11.820 1791 -
val cogvideox 1865 11.789 1788 -
val easyanimate 1801 11.384 1725 -
val pika1 1454 9.191 1393 -
val wan2 898 5.676 859 -
val kling2 838 5.297 802 -
val sora 804 5.082 771 -
Test set
test unknown 2576 21.174 2467 -
test realmotion2 2531 20.804 $24$ 11/2024
test kling1 1106 9.091 1060 2024
test hailuo 949 7.800 910 2024
test seedance 928 7.628 888 6/2025
test mochi1 580 4.767 555 9/2025
test jimeng 501 4.118 480 2025
test gen3 500 4.110 479 6/2024
test luma 500 4.110 478 4/2025
test vidu 491 4.036 469 5/2024
test pyramid 488 4.011 468 3/2025
test skyreels 399 3.280 382 4/2025
test pixverse 277 2.277 265 2025
test pika2 186 1.529 178 2025
test gen4 154 1 147 3/2025
@@ 0.B.2 Video frames from generators
Appendix 0.C Deatil config & Hardware
Table 6: Base training config detail
Category Setting Value
Optimizer Optimizer AdamW
Learning rate 3e-4
Weight decay 0.05
LR schedule Cosine decay
Warmup 1 epoch
Min learning rate 1e-6
Gradient clipping 1.0
Mixed precision AMP enabled
Random seeds 3
Model (AE-Swin base) Hidden dimension 512
Attention heads 8
Vision encoder V-JEPA2
Temporal window size 4
Spatial window size 4
Temporal blocks 2
Spatial blocks 2
Input processing Per video embeddings 16
Raw frames to V-JEPA2 32 (2-frame tubelets)
Hardware GPU RTX 6000 Ada (48GB)
VRAM used est. 42GB
Training setup Single GPU
Disk Space Video 355GB
Per embedding file (.pt) 8.1MB
Embedding total 1.1TB
Vision Encoder V-JEPA 2 vjepa2-vitl-fpc64-256
CLIP clip-vit-large-patch14
DINO-v3 dinov3-vitl16-pretrain
DINO-v2 dinov2-base
ConvNeXt convnextv2-large-22k-384
Appendix 0.D Extended results
Below, we present more experiment result. First, the hyperparameters sweep is shown in Tables 7 & 8. Second, we present the result of ablation study experimented on each vision encoders: V-JEPA 2 (Tables 9 & 10), CLIP (Tables 12 & 12), DINO-v3 (Tables 13 & 14), DINO-v2 (Tables 15 & 16).